# Group Decision-Making among Privacy-Aware Agents

### 15 Feb 2024 v2

@@ 3.4. Online Learning from Intermittent Streams
$T=\mathcal{O}\bigg{(}\sqrt{\max\bigg{\{}\frac{n|\Theta|^{2}(nQ_{n,\Theta}+V_{n,\Theta})(\max_{i\in[n]}\nu_{i}+\sqrt{\Xi_{n}/\eta})}{2l_{n,\Theta}\eta(1-(\beta_{n}^{\star})^{2})},\sqrt{\frac{n\varrho}{l_{n,\Theta}}}\bigg{\}}}\bigg{)}$
iterations, yields $\mathbb{P}\left[\Theta^{\star}\subseteq\hat{\Theta}^{\mathrm{OL}}_{i,T}\right]\geq 1-2\eta$. Subsequently, the distributions that optimize the runtime are $\mathcal{D}_{i,t}^{\star}=\mathrm{Lap}\left(\frac{\Delta_{i,t}|\Theta|}{{\varepsilon}}\right)$.
Proof Sketch. In this case, we can similarly prove that the dynamics in the log-belief ratio space are linear. The proof starts by showing that the deviation of $(1/t)\sum_{\tau=0}^{t-1}n_{i,\tau}$ from $\nu_{i}$ is no more than $\sqrt{\Xi_{n}/(\eta t)}$ with probability at least $1-\eta$. Subsequently, conditioned on this “good” event, we show that the distance between the average log-belief ratio $(1/t)\log\left(\bm{\nu}_{i,t}(\hat{\theta})/\bm{\nu}_{i,t}(\check{\theta})\right)$ between any pair of states converges to $(1/n)\sum_{i=1}^{n}\nu_{i}D_{KL}\left(\ell_{i}(\cdot|\hat{\theta})|\ell_{i}(\cdot|\check{\theta})\right)$ is inversely proportional to $\sqrt{t}$, $1-(\beta^{\star}_{n})^{2}$, and is proportional to the standard deviation of the KL divergence $Q_{n,\Theta}$ as well as $V_{n,\Theta}$. Similarly to Theorem 3.2, we threshold the log-belief ratio between an MLE state and a non-MLE state and obtain the convergence time $T$. The optimal distributions are derived similarly to Theorem 3.2. ∎
@@ 5. Discussion and Conclusions
First, for the AM/GM estimators, we conclude that introducing privacy in the distributed MLE task incurs a $\mathrm{polylog}\left(1/{\varepsilon},1/\eta\right)$ cost compared to the non-private benchmark with respect to ${\varepsilon}$ and $\eta$, as well as a $\mathrm{poly}\left(|\Theta|,\log(\Delta_{n,\Theta})\right)$ cost compared to the non-private benchmark with respect to $|\Theta|$. Second, in the two-threshold regime, the distributed MLE task incurs a $\mathrm{polylog}\left(1/{\varepsilon},1/\eta,1/q\right)$ cost with respect to ${\varepsilon}$ and $\eta$, as well as a $\mathrm{poly}\left(1/\pi,|\Theta|,\Delta_{n,\Theta},\log(1/q)\right)$ overhead with respect to $|\Theta|$.
We observe the following trade-off between the AM/GM algorithm (Algorithm 3) and the two-threshold algorithm (Algorithm 4): while the two-threshold algorithm has more free parameters and is therefore more flexible than the AM/GM algorithm, the added flexibility of the two-threshold algorithm comes at a cost to its communication complexity.
@@ A.2. Auxiliary Lemmas
guarantees that $\mathbb{E}\left[\bm{N}_{i,T}(\theta^{\star}\right]=\mathbb{P}\left[\bm{\nu}_{i,k,T}(\theta^{\star})>\tau\right]\geq q_{2}$ for all $\theta^{\star}\in\Theta^{\star}$. Then by following the analysis similar to Theorem 3.5 by utilizing the union bound and the Chernoff bound, we can prove that $\mathbb{P}[\Theta^{\star}\not\subseteq\Theta^{\mathrm{thres},2}_{i,T}]\leq|\Theta^{\star}|e^{-2K\pi_{2}^{2}}$. Subsequently, setting $K\geq\frac{\log(|\Theta^{\star}|/\eta_{2})}{2\pi_{2}^{2}}$ makes the error at most $\eta_{2}$.
@@
guarantees that $\mathbb{E}\left[\bm{N}_{i,T}(\bar{\theta})\right]=\mathbb{P}\left[\bm{\nu}_{i,k,T}(\bar{\theta})\leq\tau\right]\leq q_{1}$ for all $\bar{\theta}\in\bar{\Theta}$. Then by following the analysis similar to Theorem 3.5 by utilizing the union bound and the Chernoff bound, we can prove that $\mathbb{P}[\Theta^{\mathrm{thres},1}_{i,T}\subseteq\Theta^{\star}]\leq|\bar{\Theta}|e^{-2K\pi_{1}^{2}}$. Subsequently setting $K\geq\frac{\log(|\Theta^{\star}|/\eta_{1})}{2\pi_{1}^{2}}$ makes the error at most $\eta_{1}$.
@@ B.2. Auxiliary Lemmas
$V_{n,\Theta}^{\prime}=\max_{\bar{\theta}\in\bar{\Theta},\theta^{\star}\in\Theta^{\star}}\sqrt{\mathbb{V}\left[\frac{1}{nt}\sum_{i=1}^{n}\sum_{\tau=0}^{t-1}n_{i,\tau}\bm{\zeta}_{i,\tau}(\hat{\theta},\check{\theta})\bigg{|}B_{t}^{c}\right]}\leq\frac{\sqrt{2}}{n}(nQ_{n,\Theta}+V_{n,\Theta})\left(\max_{i\in[n]}\nu_{i}+\sqrt{\frac{\Xi_{n}}{\eta t}}\right).$
@@
Therefore, we can show that by applying a union bound over $\Theta^{\star}$, we have that with probability $1-\eta$, for all $\theta^{\star}\in\Theta$ and $\bar{\theta}\in\bar{\Theta}$
$\bm{\psi}_{i,t}(\theta^{\star},\bar{\theta})\geq\frac{t}{n}l_{n,\Theta}-\frac{|\Theta|^{2}V_{n,\Theta}^{\prime}\sqrt{t}}{2\eta(1-(\beta_{n}^{\star})^{2})}.$
To make the RHS at least the log-belief threshold $\varrho$, we require $t\geq T$ where we can calculate $T$ to be
$T=\mathcal{O}\left(\sqrt{\frac{n}{l_{n,\Theta}}\max\left\{\frac{|\Theta|^{2}(nQ_{n,\Theta}+V_{n,\Theta})(\max_{i\in[n]}\nu_{i}+\sqrt{\Xi_{n}/\eta})}{2\eta(1-(\beta_{n}^{\star})^{2})},\sqrt{\frac{l_{n,\Theta}\varrho}{n}}\right\}}\right).$
Therefore, by setting the belief threshold as a $\tau=1/(1+e^{\varrho}$ as in Theorem 3.2, we get that
$\mathbb{P}\left[\Theta^{\star}\subseteq\hat{\Theta}^{\mathrm{OL}}_{i,T}\right]\geq\mathbb{P}[B_{n,T}^{c}]\mathbb{P}\left[\Theta^{\star}\subseteq\hat{\Theta}^{\mathrm{OL}}_{i,T}|B_{n,T}^{c}\right]\geq(1-\eta)^{2}\geq 1-2\eta.$

### 12 Mar 2024 v3

@@ 3.4. Online Learning from Intermittent Streams
$T=\mathcal{O}\bigg{(}\sqrt{\max\bigg{\{}\frac{n|\Theta|(nQ_{n,\Theta}+V_{n,\Theta})(\max_{i\in[n]}\nu_{i}+\sqrt{\Xi_{n}/\eta})}{l_{n,\Theta}\eta(1-(\beta_{n}^{\star})^{2})},\sqrt{\frac{n\log(|\Theta|)}{l_{n,\Theta}}}\bigg{\}}}\bigg{)}$
iterations, yields $\mathbb{P}\left[\theta^{\star}=\hat{\theta}^{\mathrm{OL}}_{i,T}\right]\geq 1-\eta$ as long as $\sum_{i=1}^{n}\nu_{i}D_{KL}\left(\ell_{i}(\cdot|\bar{\theta})|\ell_{i}(\cdot|\theta^{\star})\right)<0$ for all $\bar{\theta}\neq\theta^{\star}$. Subsequently, the distributions that optimize the runtime are $\mathcal{D}_{i,t}^{\star}=\mathrm{Lap}\left({\Delta_{i,t}|\Theta|}/{{\varepsilon}}\right)$.
Proof Sketch. In this case, we can similarly prove that the dynamics in the log-belief ratio space are linear. The proof starts by showing that the deviation of $(1/t)\sum_{\tau=0}^{t-1}n_{i,\tau}$ from $\nu_{i}$ is no more than $\sqrt{\Xi_{n}/(\eta t)}$ with probability at least $1-\eta$. Subsequently, conditioned on this “good” event, we show that the distance between the average log-belief ratio $(1/t)\log\left(\bm{\nu}_{i,t}(\hat{\theta})/\bm{\nu}_{i,t}(\check{\theta})\right)$ between any pair of states converges to $(1/n)\sum_{i=1}^{n}\nu_{i}D_{KL}\left(\ell_{i}(\cdot|\hat{\theta})|\ell_{i}(\cdot|\check{\theta})\right)$ is inversely proportional to $\sqrt{t}$, $1-(\beta^{\star}_{n})^{2}$, and is proportional to the standard deviation of the KL divergence $Q_{n,\Theta}$ as well as $V_{n,\Theta}$. Similarly to Theorem 3.2, we threshold the log-belief ratio between $\theta^{\star}$ and any $\bar{\theta}\neq\theta^{\star}$ and obtain the convergence time $T$. The optimal distributions are derived similarly to Theorem 3.2. ∎
@@ 5. Discussion and Conclusions
First, for the AM/GM estimators, we conclude that introducing privacy in the distributed MLE task incurs a $\mathrm{polylog}\left(1/{\varepsilon},1/\eta\right)$ cost compared to the non-private benchmark with respect to ${\varepsilon}$ and $\eta$, as well as a $\mathrm{poly}\left(|\Theta|,\log(\Delta_{n,\Theta})\right)$ cost compared to the non-private benchmark with respect to $|\Theta|$. Second, in the double-threshold regime, the distributed MLE task incurs a $\mathrm{polylog}\left(1/{\varepsilon},1/\eta,1/q\right)$ cost with respect to ${\varepsilon}$ and $\eta$, as well as a $\mathrm{poly}\left(1/\pi,|\Theta|,\Delta_{n,\Theta},\log(1/q)\right)$ overhead with respect to $|\Theta|$.
We observe the following trade-off between the AM/GM algorithm (Algorithm 3) and the double-threshold algorithm (Algorithm 4): while the double-threshold algorithm has more free parameters and is therefore more flexible than the AM/GM algorithm, the added flexibility of the double-threshold algorithm comes at a cost to its communication complexity.
@@ A.2. Auxiliary Lemmas
guarantees that $\mathbb{E}\left[\bm{N}_{i,T}(\theta^{\star}\right]=\mathbb{P}\left[\bm{\nu}_{i,k,T}(\theta^{\star})>\tau\right]\geq q_{2}$ for all $\theta^{\star}\in\Theta^{\star}$. Then by following the analysis similar to Theorem 3.5 by utilizing the union bound and the Chernoff bound, we can prove that $\mathbb{P}[\Theta^{\star}\not\subseteq\hat{\Theta}^{\mathrm{thres},2}_{i,T}]\leq|\Theta^{\star}|e^{-2K\pi_{2}^{2}}$. Subsequently, setting $K\geq\frac{\log(|\Theta^{\star}|/\eta_{2})}{2\pi_{2}^{2}}$ makes the error at most $\eta_{2}$.
@@
guarantees that $\mathbb{E}\left[\bm{N}_{i,T}(\bar{\theta})\right]=\mathbb{P}\left[\bm{\nu}_{i,k,T}(\bar{\theta})\leq\tau\right]\leq q_{1}$ for all $\bar{\theta}\in\bar{\Theta}$. Then by following the analysis similar to Theorem 3.5 by utilizing the union bound and the Chernoff bound, we can prove that $\mathbb{P}[\Theta^{\mathrm{thres},1}_{i,T}\subseteq\hat{\Theta}^{\star}]\leq|\bar{\Theta}|e^{-2K\pi_{1}^{2}}$. Subsequently setting $K\geq\frac{\log(|\Theta^{\star}|/\eta_{1})}{2\pi_{1}^{2}}$ makes the error at most $\eta_{1}$.
@@ B.2. Auxiliary Lemmas
$V_{n,\Theta}^{\prime}=\max_{\bar{\theta}\in\bar{\Theta},\theta^{\star}\in\Theta^{\star}}\sqrt{\mathbb{V}\left[\frac{1}{nt}\sum_{i=1}^{n}\sum_{\tau=0}^{t-1}n_{i,\tau}\bm{\zeta}_{i,\tau}(\hat{\theta},\check{\theta})\bigg{|}B_{n,t}^{c}\right]}\leq\frac{\sqrt{2}}{n}(nQ_{n,\Theta}+V_{n,\Theta})\left(\max_{i\in[n]}\nu_{i}+\sqrt{\frac{\Xi_{n}}{\eta t}}\right).$
@@
Therefore, we can show that by applying a union bound over $\Theta^{\star}$, we have that with probability $1-\eta$, for all $\bar{\theta}\in\bar{\Theta}$
$\log\bm{\nu}_{i,t}(\bar{\theta})\leq\bm{\psi}_{i,t}(\bar{\theta},\theta^{\star})\leq-\frac{t}{n}l_{n,\Theta}+\frac{|\Theta|V_{n,\Theta}^{\prime}\sqrt{t}}{2\eta(1-(\beta_{n}^{\star})^{2})}.$
To make the RHS the log-belief threshold at most some value $-\varrho$ for some $\varrho>0$, we require
$t\geq T=\mathcal{O}\left(\sqrt{\frac{n}{l_{n,\Theta}}\max\left\{\frac{|\Theta|(nQ_{n,\Theta}+V_{n,\Theta})(\max_{i\in[n]}\nu_{i}+\sqrt{\Xi_{n}/\eta})}{\eta(1-(\beta_{n}^{\star})^{2})},\sqrt{\frac{l_{n,\Theta}\varrho}{n}}\right\}}\right).$
To determine $\varrho$, note that $\bm{\nu}_{i,t}(\bar{\theta})\leq e^{-\varrho}$ for all $\bar{\theta}\neq\theta^{\star}$, and, thus,
$\bm{\nu}_{i,t}(\theta^{\star})=1-\sum_{\bar{\theta}\neq\theta^{\star}}\bm{\nu}_{i,t}(\bar{\theta})\geq 1-(|\Theta|-1)e^{-\varrho}.$
To determine a valid value of $\varrho$ we require $1-(|\Theta|-1)e^{-\varrho}\geq e^{-\varrho}$ which yields $\varrho\geq\log(|\Theta|)$. Therefore, setting $\varrho=\log(|\Theta|)$ and subsequently
$T=\mathcal{O}\left(\sqrt{\frac{n}{l_{n,\Theta}}\max\left\{\frac{|\Theta|(nQ_{n,\Theta}+V_{n,\Theta})(\max_{i\in[n]}\nu_{i}+\sqrt{\Xi_{n}/\eta})}{\eta(1-(\beta_{n}^{\star})^{2})},\sqrt{\frac{l_{n,\Theta}\log(|\Theta|)}{n}}\right\}}\right),$
we get that
$\mathbb{P}\left[\theta^{\star}=\hat{\theta}^{\mathrm{OL}}_{i,T}\right]\geq\mathbb{P}[B_{n,T}^{c}]\mathbb{P}\left[\theta^{\star}=\hat{\theta}^{\mathrm{OL}}_{i,T}|B_{n,T}^{c}\right]\geq(1-\eta)^{2}\geq 1-2\eta.$